OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9 Answer (1, 12) (12, -1) (12, 1) (-1, -12)

OpenStudy (anonymous):

Please show work if possible

sam (.sam.):

x - 2y = 14 x=2y+14...........................(1) ------------------------------------------- x + 3y = 9 2y+14 + 3y = 9 5y+14=9 5y=-5 y=-1 ---------------------------------------------- x=2y+14 x=2(-1)+14 x=12 ===================================== x=12, y=-1